# SparQ Attention: Bandwidth-Efficient LLM Inference

### v5

@@ 6.2 End-to-End Performance
In addition to the positive microbenchmark results, we further highlight the practical improvements SparQ Attention offers by benchmarking performance of the entire Transformer model. We implemented SparQ Attention in llama.cpp, and evaluated its performance on AMD EPYC CPU systems with up to $256$GB memory. We measured the time it took to generate a single token, given an existing sequence length $S$.
The results, as seen in Figure 9, show that SparQ Attention attains speedups at all sequence lengths evaluated, compared to the dense baseline. At the longest sequence lengths considered, SparQ Attention achieves a 2.5×2.5$\times$2.5 × speedup, showcasing the benefits of reducing the data transfer associated with attention.
Figure 9: End-to-end speedup results for SparQ Attention compared to the dense baseline with Llama $2$ $7$B. This was achieved with batch size $1$ and a compression ratio of $1/8$ on CPU, with model weights represented in $8$-bits and the KV cache represented in $16$-bits. The $\mathcal{M}_{\mathrm{dense}}/\mathcal{M}_{\mathrm{SparQ}}$ line is a theoretical result, illustrating the upper bound of speedups attained with SparQ Attention.
@@ Acknowledgements
We would like to thank Daniel Justus, Paul Balança, Oscar Key and Andrew Fitzgibbon for their helpful input and feedback on this work.

### v6

@@ 6.2 End-to-End Performance
Figure 9: End-to-end CPU speedup results for SparQ Attention compared to the dense baseline with Llama $2$ $7$B. This was achieved with batch size $1$ and a compression ratio of $1/8$, with model weights represented in $8$-bits and the KV cache represented in $16$-bits. The $\mathcal{M}_{\mathrm{dense}}/\mathcal{M}_{\mathrm{SparQ}}$ line is a theoretical result, illustrating the upper bound of speedups attained with SparQ Attention.
In addition to the positive microbenchmark results, we further highlight the practical improvements SparQ Attention offers by benchmarking performance of the entire Transformer model on both CPU and GPU, implemented in llama.cpp and gpt-fast (Meta, 2023) respectively. In both cases, we measured the time it took to generate a single token, given an existing sequence length $S$.
CPU benchmarking
We evaluated CPU benchmarking performance on AMD EPYC systems with up to $256$GB memory. The results, as seen in Figure 9, show that SparQ Attention attains speedups at all sequence lengths evaluated, compared to the dense baseline. At the longest sequence lengths considered, SparQ Attention achieves a 2.5×2.5$\times$2.5 × speedup, showcasing the benefits of reducing the data transfer associated with attention.
Figure 10: End-to-end H100 speedup results for SparQ Attention compared to the dense baseline with Llama $2$ $7$B. These speedups were achieved with batch size $1$ and a compression ratio of $1/8$, with both model weights and KV cache represented in $16$ bits. The $\mathcal{M}_{\mathrm{dense}}/\mathcal{M}_{\mathrm{SparQ}}$ illustrates the theoretical maximum performance attainable under this setting.
GPU benchmarking
Our end-to-end GPU implementation is evaluated on a single H100100100100 PCIe with $80$GB memory. Despite utilising high-bandwidth memory, GPU inference still achieves end-to-end speedups when using SparQ Attention on modest sequence lengths, as seen in Figure 10.
@@ Acknowledgements
We would like to thank Oscar Key for implementing SparQ Attention on GPU and benchmarking its end-to-end performance.
In addition, we would also like to thank Daniel Justus, Paul Balança and Andrew Fitzgibbon for their helpful input and feedback on this work.